# Data scaling for navigation in unknown environments

### v1

@@ Abstract
Generalization of imitation-learned navigation policies to environments unseen in training remains a major challenge. We address this by conducting the first large-scale study of how data quantity and data diversity affect real-world generalization in end-to-end, map-free visual navigation. Using a curated 4,565-hour crowd-sourced dataset collected across 161 locations in 35 countries, we train policies for point goal navigation and evaluate their closed-loop control performance on sidewalk robots operating in four countries, covering 125 km of autonomous driving. Our results show that large-scale training data enables zero-shot navigation in unknown environments, approaching the performance of policies trained with environment-specific demonstrations. Critically, we find that data diversity is far more important than data quantity. Doubling the number of geographical locations in a training set decreases navigation errors by $\sim 15\%$, while performance benefit from adding data from existing locations saturates with very little data. We also observe that, with noisy crowd-sourced data, simple regression-based models outperform generative and sequence-based architectures. We release our policies, evaluation setup and example videos on the project page.
@@ I INTRODUCTION
We present an analysis that disentangles how training data quantity and geographic diversity individually contribute to navigation performance in unknown environments. Specifically, we focus on vision-based navigation methods that formulate navigation as an end-to-end imitation learning task, and use the number of distinct locations in the training set as a proxy measure for diversity. We train map-free point-goal [1] navigation policies using different subsets of a crowd-sourced dataset comprising $4600\text{\,}\mathrm{h}$ of demonstrations collected across 35 countries. These subsets jointly vary the number of training locations and the amount of data sampled from each location. The resulting policies’ generalization to environments not seen during training is evaluated on robots situated in four different countries. In total, our experiments comprise 125 kilometers of real-world autonomous driving, which took 120 hours to complete.
@@ III METHODS
The recorded navigation episodes do not contain information about which goal coordinate the player was trying to reach at any given time, and the routes are often circular. To sample plausible demonstrations of reaching some goal, we divide the navigation episodes into segments by finding peaks in the relationship between traveled path length and the Euclidean distance from segment start location. An example is shown in Fig. 1.
Finally, all sensors are subsampled and aligned to $4\text{\,}\mathrm{Hz}$. The player-issued robot velocity commands are resampled so that each sensor time step $t$ receives an associated $10\text{\,}\mathrm{Hz}$ chunk $a_{t}^{gt}$ of actions over the next $1\text{\,}\mathrm{s}$. Any demonstrations from locations closer than $95\text{\,}\mathrm{km}$ to our test sites are removed from the training set.
Figure 1: A crowd-sourced navigation episode segmented into 5 demonstrations for policy training.
@@
where $s_{min}$ and $s_{max}$ are the minimum and maximum scaling weights, and $|\omega|\in[0,1]$ is the largest angular velocity magnitude in the associated $\mathbf{a}_{t}^{gt}$. We set $s_{min}$ to 1 and $s_{max}$ to 10. We apply standard color jitter augmentation to the images. The $O_{t}$, $g_{t}$ and $a_{t}^{gt}$ are mirrored horizontally with 50% probability. We train all policies for two epochs, the exact number of optimizer steps depending on dataset size.
@@ IV EXPERIMENTS
Test environments. Policy performance is evaluated in 4 different locations situated in China, Kenya, Mauritius, and Botswana, demonstrated in Figure 2. The test routes represent typical navigation tasks in urban and semi-urban environments, with lengths from $250\text{\,}\mathrm{m}$ to $400\text{\,}\mathrm{m}$. Each test route is divided into multiple shorter segments $30\text{\,}\mathrm{m}$-$90\text{\,}\mathrm{m}$ in length. To complete a full test route, the robot sequentially navigates the series of segment checkpoints. In total, our test environments contain 30 individual navigation segments.
@@
The policies trained with data from a large number of different locations perform better. The zero-shot policy is capable of complex behaviors like avoiding obstacles, stopping for humans, and finding intersections. Generally, success rates of the zero-shot and in-domain versions of the large-scale policy are very close.
@@
As shown by Figure 5, the most common failure mode for the zero-shot policy is deviating from the route e.g. because of taking a wrong turn at an intersection or failing to perform a $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ turn after goal change. We hypothesize two potential causes for this behavior, the noisy robot orientation in the training data, and the absence of longer-term planning in the point goal formulation of the navigation task. The metrics in Table I provide additional insight into policy performance. Incorporating large-scale data brings benefits across all metrics except progress speed, which mostly depends on the magnitude of linear velocities that a policy commands.
@@ Results.
Figure 10: Comparison of policy architecture success rates, averaged over the segments for 3 repetitions in all test environments.
@@ V DISCUSSION & FUTURE WORK
These results suggest several promising directions for future research. Extending the analysis to other domains, such as off-road navigation or aerial robotics, would help assess the generality of our findings. We also observed variation in the utility of data across training environments, as reflected in downstream policy performance. A more fine-grained understanding of the underlying causes of this variation is an important direction for future work. Finally, while we focus on map-free navigation, many real-world systems can leverage coarse maps (e.g., OpenStreetMap) to provide additional context. Whether such information can further improve generalization of learning-based systems remains an open question. Overall, we hope these findings motivate continued investigation into dataset composition and data scaling strategies for navigation and robot learning.

### v2

@@ Abstract
Generalization of imitation-learned navigation policies to environments unseen in training remains a major challenge. We address this by conducting the first large-scale study of how data quantity and data diversity affect real-world generalization in end-to-end, map-free visual navigation. Using a curated 4,565-hour crowd-sourced dataset collected across 161 locations in 35 countries, we train policies for point goal navigation and evaluate their closed-loop control performance on sidewalk robots operating in four countries, covering 125 km of autonomous driving. Our results show that large-scale training data enables zero-shot navigation in unknown environments, approaching the performance of policies trained with environment-specific demonstrations. Critically, we find that data diversity is far more important than data quantity. Doubling the number of geographical locations in a training set decreases navigation errors by $\sim 15\%$, while performance benefit from adding data from existing locations saturates with very little data. We also observe that, with noisy crowd-sourced data, simple regression-based models outperform generative and sequence-based architectures. We release our policies, evaluation setup and example videos on the project page, lasuomela.github.io/navigation_scaling.
@@ I INTRODUCTION
We present an analysis, illustrated in Fig. LABEL:fig:pull_fig, that disentangles how training data quantity and geographic diversity individually contribute to navigation performance in unknown environments. Specifically, we focus on vision-based navigation methods that formulate navigation as an end-to-end imitation learning task, and use the number of distinct locations in the training set as a proxy measure for diversity. We train map-free point-goal [1] navigation policies using different subsets of a crowd-sourced dataset comprising $4600\text{\,}\mathrm{h}$ of demonstrations collected across 35 countries. These subsets jointly vary the number of training locations and the amount of data sampled from each location. The resulting policies’ generalization to environments not seen during training is evaluated on robots situated in four different countries. In total, our experiments comprise 125 kilometers of real-world autonomous driving, which took 120 hours to complete.
@@ III METHODS
The recorded navigation episodes do not contain information about which goal coordinate the player was trying to reach at any given time, and the routes are often circular. To sample plausible demonstrations of reaching some goal, we divide the navigation episodes into segments by finding peaks in the relationship between traveled path length and the Euclidean distance from segment start location. An example is shown in Fig. -1.
Finally, all sensors are subsampled and aligned to $4\text{\,}\mathrm{Hz}$. The player-issued robot velocity commands are resampled so that each sensor time step $t$ receives an associated $10\text{\,}\mathrm{Hz}$ chunk $\mathbf{a}_{t}^{gt}$ of actions over the next $1\text{\,}\mathrm{s}$. Any demonstrations from locations closer than $95\text{\,}\mathrm{km}$ to our test sites are removed from the training set.
Figure -1: A crowd-sourced navigation episode segmented into 5 demonstrations for policy training.
@@
where $s_{min}$ and $s_{max}$ are the minimum and maximum scaling weights, and $|\omega|\in[0,1]$ is the largest angular velocity magnitude in the associated $\mathbf{a}_{t}^{gt}$. We set $s_{min}$ to 1 and $s_{max}$ to 10. We apply standard color jitter augmentation to the images. The $\mathbf{O}_{t}$, $\mathbf{g}_{t}$ and $\mathbf{a}_{t}^{gt}$ are mirrored horizontally with 50% probability. We train all policies for two epochs, the exact number of optimizer steps depending on dataset size.
@@ IV EXPERIMENTS
Test environments. Policy performance is evaluated in 4 different locations situated in China, Kenya, Mauritius, and Botswana (Figure IV). Each test route, ranging $250\text{\,}\mathrm{m}$ to $400\text{\,}\mathrm{m}$ in length, is divided into multiple $30\text{\,}\mathrm{m}$-$90\text{\,}\mathrm{m}$ segments. To complete a full test route, the robot sequentially navigates the series of segment checkpoints. In total, our test environments contain 30 individual navigation segments. Despite appearing as simple straight-line paths in Fig. IV, in practice completing the segments requires e.g. obstacle avoidance, terrain traversability estimation, and finding correct intersections. We refer the reader to the supplemental videos for illustration of task difficulty level.
@@
The policies trained with data from a large number of different locations perform better. The zero-shot policy is capable of complex behaviors like avoiding obstacles, stopping for humans, and finding intersections. Generally, success rates of the zero-shot and in-domain versions of the large-scale policy are very close. As shown by Figure 3,
@@
the most common failure mode for the zero-shot policy is deviating from the route e.g. because of taking a wrong turn at an intersection or failing to perform a $180\text{\,}\mathrm{\SIUnitSymbolDegree}$ turn after goal change. We hypothesize two potential causes for this behavior, the noisy robot orientation in the training data, and the absence of longer-term planning in the point goal formulation of the navigation task. The metrics in Table I provide additional insight into policy performance. Incorporating large-scale data brings benefits across all metrics except progress speed, which mostly depends on the magnitude of linear velocities that a policy commands.
@@ Results.
Figure 8: Comparison of policy architecture success rates ($\uparrow$), averaged over the segments for 3 repetitions in all test environments.
@@ V DISCUSSION & FUTURE WORK
These results suggest several promising directions for future research. Extending the analysis to other domains, such as off-road navigation or aerial robotics, would help assess the generality of our findings. We also observed variation in the utility of data across training environments, as reflected in downstream policy performance. A more fine-grained understanding of the underlying causes of this variation is an important direction for future work - what kind of demonstrations are the best for learning navigation? Overall, we hope these findings motivate continued investigation into dataset composition and data scaling strategies for navigation and robot learning.
Limitations
Our experiments were conducted in only 4 different countries and environments. Verifying the findings on a set of test environments an order of magnitude larger would help confirm their validity. The study was performed in the context of PointGoal navigation, without the use of prior maps or deployment time ’memory’. This limits the difficulty level of the navigation tasks we considered - kilometer-scale segments that require planning, and exploration with dead-ends, for example, are challenging for such ’reactive’ policies.